# Sequential model confidence sets

### v1

@@ Appendix B Conditionally bounded CRPS and quantile score differences
The following results show that CRPS and quantile score differences are conditionally bounded by the predictability of the forecasts.
@@
For $F_{1},F_{2}\in\mathcal{P}(\mathbb{R})$, the CRPS difference
$\delta(y)=\textrm{CRPS}(F_{1},y)-\textrm{CRPS}(F_{2},y)$
@@
By the integral representation of the CRPS given at (21),
$\delta(y)$ $=\int\big{(}F_{1}(x)-\mathbbm{1}\{y\leq x\}\big{)}^{2}-\big{(}F_{2}(x)-\mathbbm{1}\{y\leq x\}\big{)}^{2}\,\mathrm{d}x$
$=\int F_{1}(x)^{2}-F_{2}(x)^{2}\,\mathrm{d}x+2\int_{y}^{\infty}F_{2}(x)-F_{1}(x)\,\mathrm{d}x,$
where the first summand is independent of $y\in\mathbb{R}$. The function $x\mapsto F_{2}(x)-F_{1}(x)$ is bounded in $[-1,1]$, with limit 00 as $x\to\pm\infty$ and changes sign at the crossing points of $F_{1}$ and $F_{2}$. ∎
@@
For $x_{1},x_{2}\in\mathbb{R}$ and $\tau\in(0,1)$, the quantile loss difference
$\delta(y)=\textrm{QS}_{\tau}(x_{1},y)-\textrm{QS}_{\tau}(x_{2},y)$
is bounded in $y\in\mathbb{R}$ with
$(\tau-1)\lvert x_{1}-x_{2}\rvert\leq\delta(y)\leq\tau\lvert x_{1}-x_{2}\rvert.$
@@
Assume w.l.o.g. that $x_{1}\leq x_{2}$. Then, we have
$\delta(y)$ $=\big{(}\mathbbm{1}\{y\leq x_{1}\}-\tau\big{)}(x_{1}-y)-\big{(}\mathbbm{1}\{y\leq x_{2}\}-\tau\big{)}(x_{2}-y)$
$=y\mathbbm{1}\{x_{1}<y\leq x_{2}\}-\tau(x_{1}-x_{2})+x_{1}\mathbbm{1}\{y\leq x_{1}\}-x_{2}\mathbbm{1}\{y\leq x_{2}\}$
@@
and the bounds follow directly by differentiating cases. ∎
Appendix C Sequential family-wise error rate and family-wise validity
Let $(E_{1,t})_{t\in\mathbb{N}},\dots,(E_{m,t})_{t\in\mathbb{N}}$ be e-processes for some hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}\subseteq\mathfrak{B}(\Omega)$.
@@ Appendix C Sequential family-wise error rate and family-wise validity
Following Vovk and Wang (2021), we call $f$ a symmetric e-merging function if it is invariant with respect to permutations of its arguments and if $f(E_{i},i\in I)$ is an e-value for any family of e-values $\{E_{i}\mid i\in I\}$. The most important e-merging function is the arithmetic mean as it essentially dominates any symmetric e-merging function (Vovk and Wang, 2021, Proposition 3.1.). Let $f$ be any symmetric e-merging function. Then, by the closure principle (Vovk and Wang, 2021, Appendix A.3.), for any $t\in\mathbb{N}$,
$E^{\star}_{i,t}=\min\{f(E_{j,t},j\in I)\mid I\subseteq\{1,\dots,m\},i\in I\},\quad i=1,\dots,m,$ (22)
@@
are family-wise valid e-variables for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}$. That is, there exists a dominating family of e-variables $\{E^{\mathbb{Q}}\mid\mathbb{Q}\in\mathfrak{B}(\Omega)\}$, with $E^{\mathbb{Q}}\geq E^{\star}_{i,t}$ for any $i=1,\dots,m$ and $\mathbb{Q}\in\mathcal{H}_{i}$.
@@
If we adjust the e-processes $(E_{1,t})_{t\in\mathbb{N}},\dots,(E_{m,t})_{t\in\mathbb{N}}$ by (22) with respect to the same e-merging function $f$ for all $t\in\mathbb{N}$, then it follows directly that $(E^{\star}_{1,t})_{t\in\mathbb{N}},\dots,(E^{\star}_{m,t})_{t\in\mathbb{N}}$ are family-wise valid e-processes for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}$. Importantly, the corresponding sequential test given at (8) control the family-wise error rate (FWER) uniformly in time, that is, for any $I\subseteq\{1,\dots,m\}$ and $\mathbb{Q}\in\bigcap_{i\in I}\mathcal{H}_{i}$,
$\textrm{FWER}_{I}(\mathbb{Q})=\mathbb{Q}\left(\exists t\in\mathbb{N}:E_{i,t}^{\star}\geq 1/\alpha\textrm{ for some }i\in I\right)\leq\alpha,\quad\textrm{forall }\alpha\in(0,1).$ (23)
To verify (23), consider a dominating family of e-processes $\{(E_{t}^{\mathbb{Q}})_{t\in\mathbb{N}}\mid\mathbb{Q}\in\mathfrak{B}(\Omega)\}$. Then, for any $\alpha\in(0,1)$, $I\subseteq\{1,\dots,m\}$ and $\mathbb{Q}\in\bigcap_{i\in I}\mathcal{H}_{i}$,
@@
$\textrm{FWER}_{I}(\mathbb{Q})\leq\mathbb{Q}\left(\exists t\in\mathbb{N}:E_{t}^{\mathbb{Q}}\geq 1/\alpha\right)\leq\alpha.$
@@ Appendix D SMCSs by FDR control
For $i=1,\dots,m$, let $(\Psi_{i,t})_{t\in\mathbb{N}}$ be a sequential testing procedure for some hypothesis $\mathcal{H}_{i}\subseteq\mathfrak{B}(\Omega)$. We let $R_{t}=\sum_{i=1}^{m}\Psi_{i,t}$ be the number of all rejected hypotheses (discoveries) at $t\in\mathbb{N}$. For $\mathbb{Q}\in\bigcup_{i=1}^{m}\mathcal{H}_{i}$, we let $F_{t}(\mathbb{Q})=\sum_{i\in I_{\mathbb{Q}}}\Psi_{i,t}$ be the number of all true null hypotheses that are rejected at time $t$. The false discovery rate (FDR) at $t\in\mathbb{N}$ of the given testing procedures is defined as
@@
We say that the sequential tests $(\Psi_{1,t})_{t\in\mathbb{N}},\dots,(\Psi_{m,t})_{t\in\mathbb{N}}$ time-uniformly control the false discovery rate at level $\alpha\in(0,1)$ if, for all (possibly infinite) stopping times $\tau$,
$\textrm{FDR}_{\tau}\leq\alpha.$ (24)
@@
Wang and Ramdas (2022) study FDR control with e-values in a nonsequential setting. Looking carefully at the proof of their Theorem 2, we see that their arguments directly extend to e-processes to obtain sequential tests that satisfy (24) as explained in the following.
Let $(E_{1,t})_{t\in\mathbb{N}},\dots,(E_{m,t})_{t\in\mathbb{N}}$ be e-processes for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}\subseteq\mathfrak{B}(\Omega)$. Following Wang and Ramdas (2022), we let $E_{[1],t}\geq\dots\geq E_{[m],t}$ be the order statistics of $E_{1,t},\dots,E_{m,t}$ in decreasing order. For a given $\alpha\in(0,1)$, we define
$i^{\star}_{t}=\max\left\{i\in\{1,\dots,m\}\;\Big{|}\;\frac{iE_{[i],t}}{m}\geq\frac{1}{\alpha}\right\},\quad t\in\mathbb{N}.$
Then, we have a time-uniform FDR control at level $\alpha$ for the sequential tests
$\Psi_{i,t}=(\mathbbm{1}\{E_{i,t}\geq E_{[i_{t}^{\star}],t}\}),\quad i=1,\dots,m,t\in\mathbb{N}.$ (25)
As our construction for the strong and uniformly weak hypothesis builds on e-processes, we can use the sequential tests at (25) with respect to the e-processes given at (9). That is, at each $t\in\mathbb{N}$, we reject the $i^{\star}_{t}$ hypotheses with the largest values of the corresponding e-processes, and obtain the alternative SMCS
ˇ$\check{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid E_{i\cdot,t}<E_{[i_{t}^{\star}]\cdot,t}\right\}.$
Importantly, this sequence does not necessarily satisfy the originally imposed coverage requirement at (4) or (5) anymore. In contrast, it bounds the expected fraction of wrongly rejected models by the total number of rejected models at any random time by $\alpha$, that is, for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$ and any random stopping time $\tau$, we have
@@
$\mathbb{E}_{\mathbb{Q}}\left[\frac{\lvert\check{\mathcal{M}}^{c}_{\tau}\cap\mathcal{M}^{\star}_{\tau}\rvert}{\max\{1,\lvert\check{\mathcal{M}}_{\tau}^{c}\rvert\}}\right]\leq\alpha,$
where ˇ$\check{\mathcal{M}}^{c}_{\tau}=\{1,\dots,m\}\setminus\check{\mathcal{M}}_{\tau}$ and $(\mathcal{M}^{\star})_{t\in\mathbb{N}}$ is given by (1) or (2).
Appendix E Proof of Lemma 3.5
Proof.
For $t\in\mathbb{N}$, we have $C_{t,1-\alpha}=\{\boldsymbol{X}\in\mathbb{R}_{0}^{m\times m}\mid\lVert f_{t}(\boldsymbol{X})\rVert_{1}\leq m(m-1)/\alpha\}$, for
$f_{t}:\mathbb{R}_{0}^{m\times m}\to\mathbb{R}_{0}^{m\times m},\quad\boldsymbol{X}\mapsto\left(f_{ij,t}(X_{ij})\right)_{i\neq j}=\left(\exp\left\{\lambda_{ij}t\hat{\Delta}_{ij}-\lambda_{ij}tx-\psi_{E,c_{ij}}(\lambda_{ij})V_{ij,t}\right\}\right)_{i\neq j},$
@@
and where $\lVert\boldsymbol{X}\rVert_{1}=\sum_{i,j=1}^{m}\lvert X_{ij}\rvert$ for $\boldsymbol{X}\in\mathbb{R}^{m\times m}$. For any $i\neq j$, $f_{ij,t}$ is a nonnegative and nonincreasing convex function. Thus, the second claim follows immediately since $\boldsymbol{A}\leq\boldsymbol{B}$ implies $f_{t}(\boldsymbol{A})\geq f_{t}(\boldsymbol{B})$. Let $\boldsymbol{A},\boldsymbol{B}\in C_{t,1-\alpha}$, then, for any $\lambda\in[0,1]$,
$\lVert f_{t}\big{(}\lambda\boldsymbol{A}+(1-\lambda)\boldsymbol{B}\big{)}\rVert_{1}\leq\lVert\lambda f_{t}(\boldsymbol{A})+(1-\lambda)f_{t}(\boldsymbol{B})\rVert_{1}\leq\lambda\lVert f_{t}(\boldsymbol{A})\rVert_{1}+(1-\lambda)\lVert f_{t}(\boldsymbol{B})\rVert_{1}\leq\frac{m(m-1)}{\alpha},$

### v3

@@ Appendix B Conditionally bounded score differences
The following results show that CRPS and quantile score differences are conditionally bounded due to the predictability of the forecasts.
@@
We have $\delta(y)=\int F_{1}(x)^{2}-F_{2}(x)^{2}\,\mathrm{d}x+2\int_{y}^{\infty}F_{2}(x)-F_{1}(x)\,\mathrm{d}x,$ where the first summand is independent of $y\in\mathbb{R}$. The function $x\mapsto F_{2}(x)-F_{1}(x)$ is bounded in $[-1,1]$, with limit 00 as $x\to\pm\infty$ and changes sign at the crossing points of $F_{1}$ and $F_{2}$. ∎
@@
For $x_{1},x_{2}\in\mathbb{R}$ and $\tau\in(0,1)$, the difference $\delta(y)=\textrm{QS}_{\tau}(x_{1},y)-\textrm{QS}_{\tau}(x_{2},y)$ is bounded in $y\in\mathbb{R}$ with $(\tau-1)\lvert x_{1}-x_{2}\rvert\leq\delta(y)\leq\tau\lvert x_{1}-x_{2}\rvert.$
@@
If $x_{1}\leq x_{2}$, then $\delta(y)=y\mathbbm{1}_{\{x_{1}<y\leq x_{2}\}}-\tau(x_{1}-x_{2})+x_{1}\mathbbm{1}_{\{y\leq x_{1}\}}-x_{2}\mathbbm{1}_{\{y\leq x_{2}\}}$ and the bounds follow directly by differentiating cases. The case $x_{1}>x_{2}$ is analogous. ∎
Appendix C Multiple testing and proof of Theorem 3.1
We give some preliminaries on multiple testing corrections with e-processes before stating a proof for Theorem 3.1.
@@
Following Vovk and Wang (2021), we call $f$ a symmetric e-merging function if it is invariant with respect to permutations of its arguments and if $f(E_{i},i\in I)$ is an e-value for any family of e-values $\{E_{i}\mid i\in I\}$ and any finite index set $I$. The most important e-merging function is the arithmetic mean as it essentially dominates any symmetric e-merging function (Vovk and Wang, 2021, Proposition 3.1.). The function $f$ is defined on $\bigcup_{i=1}^{\infty}\mathbb{R}^{i}$. In other words, it can be seen as a collection of functions on $\mathbb{R}^{i}$ for each $i\in\mathbb{N}$.
@@
In the spirit of Vovk and Wang (2021), we say that e-processes $(E_{1,t}^{\star})_{t\in\mathbb{N}},\dots,(E_{m,t}^{\star})_{t\in\mathbb{N}}$ for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}$ are family-wise valid (FWV) if there exists a dominating family of e-processes, that is, if there exists a family $\{(E^{\mathbb{Q}}_{t})_{t\in\mathbb{N}}\mid\mathbb{Q}\in\mathfrak{B}(\Omega)\}$ such that $(E^{\mathbb{Q}}_{t})_{t\in\mathbb{N}}$ is an e-process for ${}\{\mathbb{Q}\}$ for all $\mathbb{Q}\in\mathfrak{B}(\Omega)$, and for any $i=1,\dots,m$ with $\mathbb{Q}\in\mathcal{H}_{i}$, it holds that $E_{t}^{\mathbb{Q}}\geq E_{i,t}^{\star}$.
Let $(E_{1,t})_{t\in\mathbb{N}},\dots,(E_{m,t})_{t\in\mathbb{N}}$ be e-processes for some hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}\subseteq\mathfrak{B}(\Omega)$, respectively, and let $f$ be a symmetric e-merging function. Then, by the closure principle (Marcus et al., 1976), for any $t\in\mathbb{N}$, $E^{\star}_{i,t}=\min\{f(E_{j,t},j\in I)\mid I\subseteq\{1,\dots,m\},i\in I\},i=1,\dots,m,$ are family-wise valid e-variables for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}$, see Vovk and Wang (2021, Section 5). That is, there exists a dominating family of e-variables $\{E_{t}^{\mathbb{Q}}\mid\mathbb{Q}\in\mathfrak{B}(\Omega)\}$, with $E_{t}^{\mathbb{Q}}\geq E^{\star}_{i,t}$ for any $i=1,\dots,m$ and $\mathbb{Q}\in\mathcal{H}_{i}$. Indeed, $E_{t}^{\mathbb{Q}}$ is explicitly given by $E_{t}^{\mathbb{Q}}=f(E_{i,t},i\in I_{\mathbb{Q}})$, where $I_{\mathbb{Q}}=\{i\mid\mathbb{Q}\in\mathcal{H}_{i}\}$. (If $I_{\mathbb{Q}}=\emptyset$, then $E_{t}^{\mathbb{Q}}=1$.)
@@
Suppose that we adjust the e-processes $(E_{1,t})_{t\in\mathbb{N}},\dots,(E_{m,t})_{t\in\mathbb{N}}$ by the closure principle with respect to the same e-merging function $f$ for all $t\in\mathbb{N}$. Then it follows that $(E^{\star}_{1,t})_{t\in\mathbb{N}},\dots,(E^{\star}_{m,t})_{t\in\mathbb{N}}$ are family-wise valid e-processes for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}$, since $(E_{t}^{\mathbb{Q}})_{t\in\mathbb{N}}$ are e-processes. To see this, let $\tau$ be a stopping time. Then, $\mathbb{E}_{\mathbb{Q}}(E_{\tau}^{\mathbb{Q}})=\mathbb{E}_{\mathbb{Q}}(f(E_{i,\tau},i\in I_{\mathbb{Q}}))\leq 1$, since $E_{i,\tau}$, $i\in I_{\mathbb{Q}}$ are e-values under $\mathbb{Q}$.
Proof of Theorem 3.1.
Let $\mathbb{Q}\in\mathfrak{B}(\Omega)$. We have $i\in\mathcal{M}^{\bullet,\star}$ if and only if $\mathbb{Q}\in\mathcal{H}^{\bullet}_{i\cdot}$. Therefore,
$\mathbb{Q}(\exists t\in\mathbb{N}:\mathcal{M}^{\bullet,\star}\not\subseteq\widehat{\mathcal{M}}_{t})$ $=\mathbb{Q}(\exists t\in\mathbb{N}:E_{i\cdot,t}^{\star}\geq 1/\alpha\textrm{ for some }i\in\mathcal{M}^{\bullet,\star})$
$=\mathbb{Q}\left(\bigcup_{t\in\mathbb{N}}\bigcup_{i\in\mathcal{M}^{\bullet,\star}}\{E_{i\cdot,t}^{\star}\geq 1/\alpha\}\right)$
@@
$\leq\mathbb{Q}\left(\bigcup_{t\in\mathbb{N}}\bigcup_{i\in\mathcal{M}^{\bullet,\star}}\{E_{t}^{\mathbb{Q}}\geq 1/\alpha\}\right)=\mathbb{Q}(\exists t\in\mathbb{N}:E_{t}^{\mathbb{Q}}\geq 1/\alpha)\leq\alpha,$
where $E_{t}^{\mathbb{Q}}=f(E_{i\cdot,t},i\in I_{\mathbb{Q}})$ with $I_{\mathbb{Q}}=\{i\mid\mathbb{Q}\in\mathcal{H}^{\bullet}_{i\cdot}\}$. It follows directly that the running intersection ~$\widetilde{\mathcal{M}}_{t}=\bigcap_{r\leq t}\widehat{\mathcal{M}}_{r},t\in\mathbb{N}$, is an SMCS as well.∎
Remark 3.
For the construction of SMCSs in Section 3.2, we merged the pairwise e-processes $(E_{ij,t})_{t\in\mathbb{N}}$ into a single e-process $E_{i\cdot,t}^{\star}$ by averaging over $j$, and then adjusted the merged e-processes using the closure principle. Alternatively, we could directly work with the pairwise e-processes without first merging them. More specifically, starting with the pairwise e-processes $(E_{ij,t})_{t\in\mathbb{N}}$, for all pairs $(i,j)\in\mathcal{M}_{0}$, one can directly use the closure principle (Vovk and Wang, 2021, Algorithm 1) to derive the adjusted e-processes
$E^{\star}_{ij,t}=\min_{I\subseteq\{1,\dots,m\}^{2}:(i,j)\in I}\frac{1}{\lvert I\rvert}\sum_{(k,\ell)\in I}E_{k\ell,t},\quad i,j\in\mathcal{M}_{0},t\in\mathbb{N}.$
Finally, for some significance level $\alpha\in(0,1)$, we would define
^$\widehat{\mathcal{M}}_{t}=\left\{i\in\mathcal{M}_{0}\mid E_{ij,t}^{\star}<1/\alpha,\text{ for all }j\in\mathcal{M}_{0}\right\},\quad t\in\mathbb{N}.$
However, the corresponding tests would be less powerful, since for any $i\in\mathcal{M}_{0}$ and $j\neq i$,
$\allowdisplaybreaks E^{\star}_{ij,t}$ $=\min\left\{\frac{1}{\lvert I\rvert}\sum_{(k,l)\in I}E_{kl,t}\;\Big{|}\;I\subseteq\{1,\dots,m\}^{2},(i,j)\in I\right\}$
@@ Appendix D SMCSs by FDR control
For $i=1,\dots,m$, let $(\Psi_{i,t})_{t\in\mathbb{N}}$ be a sequential test for some hypothesis $\mathcal{H}_{i}\subseteq\mathfrak{B}(\Omega)$. We let $R_{t}=\sum_{i=1}^{m}\Psi_{i,t}$ be the number of all rejected hypotheses (discoveries) at $t\in\mathbb{N}$. For $\mathbb{Q}\in\bigcup_{i=1}^{m}\mathcal{H}_{i}$, we let $F_{t}(\mathbb{Q})=\sum_{i\in I_{\mathbb{Q}}}\Psi_{i,t}$ be the number of all true null hypotheses that are rejected at time $t$. The false discovery rate (FDR) at $t\in\mathbb{N}$ of the given testing procedures is defined as
@@
We say that the sequential tests $(\Psi_{1,t})_{t\in\mathbb{N}},\dots,(\Psi_{m,t})_{t\in\mathbb{N}}$ time-uniformly control the false discovery rate at level $\alpha\in(0,1)$ if, for all (possibly infinite) stopping times $\tau$, $\textrm{FDR}_{\tau}\leq\alpha.$ Wang and Ramdas (2022) study FDR control with e-values in a nonsequential setting. Looking carefully at the proof of their Theorem 2, we see that their arguments directly extend to e-processes to obtain sequential tests that time-uniformly control FDR as explained in the following.
Let $(E_{1,t})_{t\in\mathbb{N}},\dots,(E_{m,t})_{t\in\mathbb{N}}$ be e-processes for the hypotheses $\mathcal{H}_{1},\dots,\mathcal{H}_{m}\subseteq\mathfrak{B}(\Omega)$. Following Wang and Ramdas (2022), we let $E_{[1],t}\geq\dots\geq E_{[m],t}$ be the order statistics of $E_{1,t},\dots,E_{m,t}$ in decreasing order and define $i^{\star}_{t}=\max\{i\in\{1,\dots,m\}\;|\;{iE_{[i],t}}\geq m/{\alpha}\}$, for a given $\alpha\in(0,1)$. Then, we have a time-uniform FDR control at level $\alpha$ for the sequential tests $\Psi_{i,t}=(\mathbbm{1}\{E_{i,t}\geq E_{[i_{t}^{\star}],t}\}),i=1,\dots,m,t\in\mathbb{N}.$ As our construction for the strong and uniformly weak hypothesis builds on e-processes, we can use the above sequential tests with respect to the merged e-processes $(E_{1\cdot,t})_{t\in\mathbb{N}},\dots,(E_{m\cdot,t})_{t\in\mathbb{N}}$ given in Section 3.2. That is, at each $t\in\mathbb{N}$, we reject the $i^{\star}_{t}$ hypotheses with the largest values of the corresponding e-processes, and obtain the alternative SMCS ˇ$\check{\mathcal{M}}_{t}=\{i\in\mathcal{M}_{0}\mid E_{i\cdot,t}<E_{[i_{t}^{\star}]\cdot,t}\}$. Importantly, this sequence does not necessarily satisfy the time-uniform coverage property from Section 2 anymore. In contrast, it bounds the expected ratio of wrongly rejected models to the total number of rejected models at any random time by $\alpha$, that is, for any $\mathbb{Q}\in\mathfrak{B}(\Omega)$ and any random stopping time $\tau$, we have $\mathbb{E}_{\mathbb{Q}}\big{[}{\lvert\check{\mathcal{M}}^{c}_{\tau}\cap\mathcal{M}^{\star}_{\tau}\rvert/\max\{1,\lvert\check{\mathcal{M}}_{\tau}^{c}\rvert\}}\big{]}\leq\alpha,$ where ˇ$\check{\mathcal{M}}^{c}_{\tau}=\{1,\dots,m\}\setminus\check{\mathcal{M}}_{\tau}$ and $(\mathcal{M}^{\star})_{t\in\mathbb{N}}$ is given by (1) or (2).
Appendix E Proofs for Section 3.3
Proof of Theorem 3.4.
For any $\mathbb{Q}\in\mathfrak{B}(\Omega)$, we have $\mathbb{Q}(\exists t\in\mathbb{N}:\mathcal{M}^{\textrm{w},\star}_{t}\not\subseteq\widehat{\mathcal{M}}_{t})\leq\mathbb{Q}(\exists t\in\mathbb{N}:\boldsymbol{\Delta}_{t}\notin C_{t})\leq\alpha$ by Ville’s inequality. Moreover, we have
$\mathbb{Q}(\forall t\geq 1:\mathcal{M}^{\textrm{uw},\star}\subseteq\widetilde{\mathcal{M}}_{t})$ $=\mathbb{Q}\left(\forall t\geq 1:\mathcal{M}^{\textrm{uw},\star}\subseteq\widehat{\mathcal{M}}_{t}\right)$
@@
$\geq\mathbb{Q}\left(\forall t\geq 1:\mathcal{M}^{\textrm{w},\star}_{t}\subseteq\widehat{\mathcal{M}}_{t}\right)\geq 1-\alpha.$
Proof of Proposition 3.5.
@@
For $t\in\mathbb{N}$, we have $C_{t,1-\alpha}=\{\boldsymbol{X}\in\mathbb{R}_{0}^{m\times m}\mid\lVert M_{t}(\boldsymbol{X})\rVert_{1}\leq m(m-1)/\alpha\}$, for $M_{t}:\mathbb{R}_{0}^{m\times m}\to\mathbb{R}_{0}^{m\times m},\boldsymbol{X}\mapsto\left(M_{ij,t}(X_{ij})\right)_{i\neq j}$ where $\lVert\boldsymbol{X}\rVert_{1}=\sum_{i,j=1}^{m}\lvert X_{ij}\rvert$ for $\boldsymbol{X}\in\mathbb{R}^{m\times m}$. Recall that, for any $i\neq j$, $M_{ij,t}$ is a nonnegative, convex and decreasing function. Thus, the second claim follows immediately since $\boldsymbol{A}\leq\boldsymbol{B}$ implies $M_{t}(\boldsymbol{A})\geq M_{t}(\boldsymbol{B})$. Let $\boldsymbol{A},\boldsymbol{B}\in C_{t,1-\alpha}$, then, for any $\lambda\in[0,1]$,
$\lVert M_{t}\big{(}\lambda\boldsymbol{A}+(1-\lambda)\boldsymbol{B}\big{)}\rVert_{1}$ $\leq\lVert\lambda M_{t}(\boldsymbol{A})+(1-\lambda)M_{t}(\boldsymbol{B})\rVert_{1}$
$\leq\lambda\lVert M_{t}(\boldsymbol{A})\rVert_{1}+(1-\lambda)\lVert M_{t}(\boldsymbol{B})\rVert_{1}\leq\frac{m(m-1)}{\alpha},$